This problem has me confused. I have tried using the Ztest and even Chi Sq but cannot determine how to compute using 2 standard deviations. Thanks in advance.
A lab technician is tested for her consistency by taking multiple measurements of cholesterol levels from the same blood sample. The target accuracy is a standard deviation of 1.2 or less. If 20 measurements are taken and their standard deviation is 1.8, is there enough evidence to reject the claim that her standard deviation is within the limit at  a= .01?
A) No, comparing the test value 45.00 and the critical value 37.566.
B) Yes, comparing the test value 42.75 and the critical value 36.191.
C) Yes, comparing the test value 45.00 and the critical value 37.566.
D) No, comparing the test value 42.75 and the critical value 36.191.

Answer by stanbon(75887)

A lab technician is tested for her consistency by taking multiple measurements of cholesterol levels from the same blood sample. The target accuracy is a standard deviation of 1.2 or less. If 20 measurements are taken and their standard deviation is 1.8, is there enough evidence to reject the claim that her standard deviation is within the limit at a= .01?
----------------------------------
Ho: signa >= 1.2
Ha: sigma < 1.2 (claim)
---------
alpha = 1%
Critical Value: Chi-Sq = 36.191
test stat: Chi-Sq = (n-1)s^2/sigma^2 = 19*1.8^2/1.2^2 = 42.75
--
Conclusion: Since the test stat is in the reject interval,
reject Ho.
The test results support the claim.
=================
Ans: B
=================
Cheers,
Stan H.




A) No, comparing the test value 45.00 and the critical value 37.566.
B) Yes, comparing the test value 42.75 and the critical value 36.191.
C) Yes, comparing the test value 45.00 and the critical value 37.566.
D) No, comparing the test value 42.75 and the critical value 36.191.